# Infinigen-Sim: Procedural Generation of Articulated Simulation Assets

### v2

@@ Abstract
We introduce Infinigen-Sim, a toolkit which enables users to create diverse and realistic articulated object procedural generators. These tools are composed of high-level utilities for use creating articulated assets in Blender, as well as an export pipeline to integrate the resulting assets into common robotics simulators. We demonstrate our system by creating procedural generators for 5 common articulated object categories. Experiments show that assets sampled from these generators are useful for movable object segmentation, training generalizable reinforcement learning policies, and sim-to-real transfer of imitation learning policies.
Keywords: Robot Simulation, Articulated Objects, Procedural Generation
1 Introduction
Figure 1: Panels (a)-(e) show sampled assets generated using Infinigen-Sim. We develop generators for the door, toaster, refrigerator, dishwasher, and lamp categories. As shown in Panel (f), our assets are can be directly integrated within Infinigen [1] scenes to train robot policies in highly diverse, photorealistic, large-scale simulation environments.
@@ 1 Introduction
Interacting with articulated objects is an essential step in many important applications for robotics. From using hinged fridges and dishwashers, pushing buttons, opening drawers, to traversing doors with various handles, robots require the ability to manipulate articulated objects to accomplish useful tasks in the real world.
@@
Learning in physical simulation [2, 3, 4, 5, 6, 7] has proven a useful strategy for many robotics tasks [8, 9, 10, 11, 12, 13, 14, 9]. To best use this strategy for articulated object manipulation, we require large-scale datasets of articulated assets with as great a diversity of geometry and joints as possible. Articulated assets are challenging to acquire, as objects curated from the internet [15, 16] or 3D-scanning [17] are typically static and lack joint annotations. Widely used articulated object datasets instead use labor-intensive human annotation [18, 19, 20, 21, 22] for joint position and axes, which limits the number of unique geometries and articulation annotations.
To this end, we propose Infinigen-Sim, a tool which enables the creation of diverse articulated assets through procedural generation. To complement existing sources of data, we envision that users can choose an object class of interest (e.g. lamps, doors) and create a procedural generator for it, which is a 3D graphics program that outputs assets of the class as a function of optionally randomized parameters (e.g. dimensions, part placement). Our approach has many advantages:
@@
Unlimited kinematic variation: Procedural generators can sample assets with dense coverage of important physical dimensions, e.g. the handle-to-hinge distance of a door. This means interacting with the assets requires different trajectories, e.g. different radius arcs of an end effector opening doors with different handle-to-hinge distances. Continuous variation also applies to geometry detail (e.g. edge bevel radii) and tolerances (e.g. toaster slot width), which all serve to create dense coverage and diverse training data.
@@
Combinatorial coverage: Procedural generators can randomly swap sub-parts to use different procedural generators (e.g. drastic changes in handle style: knobs, levers, crash-bars). We can also easily express variable-quantity arrays of articulated parts (e.g. 1-5 fridge drawers), as well as very small parts (e.g. tiny buttons on a toaster), which be challenging for annotators or image-based capture.
Uniformly high-quality: Each procedural generator is code that can be verified for correctness. This means our joint locations and axes do not make errors for individual asset instances as can happen for human annotation. Our procedural assets also have realistic materials (derived from Infinigen-Indoors) and detailed but efficient geometry with no holes or protruding triangles.
@@
Infinigen-Sim builds on 3D-graphics and procedural generation tools from Blender [23] and Infinigen Indoors [1], but adds tools and generators designed for articulated object generation. Specifically, we augment Blender’s Geometry-Nodes system with new modules (known as node-groups) which implement hinged and sliding joints for procedural assets and ensure that these joints are represented in a unified format for use by downstream tools. We provide export code which converts assets created using these tools into high quality assets for simulation in popular formats such as URDF, USD, and MJCF.
Experiments show that objects created using procedural generators created with Infinigen-Sim are useful for tasks in both vision and robot learning. Specifically, we used Infinigen-Sim to create articulated procedural generators for 5 object categories, namely: doors, toasters, fridges, dishwashers and lamps (shown in Fig. 1). We use these to generate datasets for movable part segmentation and achieve better generalization for more detailed articulated parts. We also train both RL and imitation learning policies on simulation data and show improved generalization to unseen object instances and demonstrated successful sim-to-real transfer through real-world experiments on a Franka Panda robot arm.
We hope Infinigen-Sim will complement existing articulated assets by providing dense high-quality coverage for specific object classes, and we hope this library of pre-implemented object generators will grow through open-source engagement with the community.
2 Related Work
@@
Asset Datasets for Robot Simulation. Although works have collected large-scale static 3D asset datasets [16, 15], interactive and articulated 3D datasets [18, 24, 25, 26, 22] remain limited in both quantity and quality. Sapien [18] collected 2k+ assets from PartNet [24] and manually annotated the articulated joints. Others [19, 20, 27] scanned real-world objects to curate separate articulated object datasets. However, issues like inaccurate joint annotations and a lack of kinematic diversity remain. Infinigen-Sim instead adopts a new approach (procedural generation) which has precise ground truth and is highly scalable, only requiring a new procedural generator for each object class, with no human time cost for each unique object geometry or articulation configuration.
@@
Articulated Object Generation. Recent work has exploited AI techniques to augment and diversify static asset datasets [16, 28] to create articulated objects [29, 30]. However, these systems lack controllability of asset geometries and do not guarantee correct joint configurations [31]. Prior work [32, 29, 33, 34, 35, 36, 30, 37] has also generated articulated assets with real-to-sim pipelines. Paris [34] learns part-based neural radiance fields to reconstruct articulated assets. Real2Code [33] and URDFormer [35] constructs URDFs with RGB images as input, using a mixture of off-the-shelf vision models and LLMs. That being said, real-to-sim approaches often fail to construct articulated assets with high-quality. Additionally, these methods fall short on capturing occluded articulations and tend to produce inaccurate geometries and joint annotations [31]. Inherited from Infinigen-Indoors [1], our tool allows users to model each individual part from scratch including its joint properties. This means users can also easily create high quality articulated bodies that may be occluded by outer surfaces (see Fig. 1(c) and (d)).
@@
Procedural Generation. Many prior works create procedural training data focused on indoor environments [5, 38], nature [38] and cities [39] but do not create procedural articulated objects. More similar to our work is Eppner et al. [40], which introduces a scene generator including some procedural articulated assets made in Python. However, their assets are composed mainly of mesh primitives (cubes/cylinders), do not contain multi-level articulations (e.g. pivoting handle on a pivoting door), and is extensible only by programmers. Our system is constructed using Blender, has detailed geometry and materials, and extensible by both programmers and anyone familiar with Blender.
3 Infinigen-Sim
Infinigen-Sim is designed to aid in creating procedural generators which produce simulation-ready articulated assets. It is composed of a set of procedural articulation tools, followed by export code which produces a file for use in robotics simulators. This overall workflow is visualized in Fig. 2.
@@ 3 Infinigen-Sim
Figure 2: Users create a procedural asset using Blender’s Geometry Nodes feature and our custom joint nodes. The node graph is then converted to Python code and a corresponding blueprint storing the kinematic structure is created. We sample parameters to generate the asset and pass both the asset and blueprint to the exporter, which recursively builds a simulation-ready articulated object.
3.1 Articulation Procedural Generator Tools
@@
Infinigen-Sim leverages Blender’s procedural modeling tools [23], especially Geometry Nodes and Shader Nodes, which are an artist-friendly, GUI-based, domain specific language for procedural generation. These tools allow users to create meshes and materials by composing primitives, geometric transformations, and scalar vector arithmetic, all represented as nodes in a direct acyclic graph. Geometry nodes are widely adopted among 3D artists in the modeling community, with extensive documentation and a large ecosystem of tutorials and resources. These node-based tools are general purpose and can represent essentially all desirable object geometries.
@@
To design articulated assets within this system, we create two new nodes that implement revolute and prismatic joints. These nodes accept two incoming geometries (defined via prior nodes) and introduce a parent-child relationship annotated with the relevant joint type. Each joint node also allows users to set the pivot position, axis, and joint range, either as fixed values or through parameters passed from other nodes. Users can leverage Blender’s built-in math nodes to compute these values directly from the asset’s geometry within the node graph. Several example uses of these joint nodes are shown in Appendix Sec. A.2.
@@
These joint nodes have three purposes. First, they directly speed up the procedural generator creation process, since they implement a commonly used operation which no longer needs to be re-implemented for each object that uses a hinge or sliding joint. Second, they provide immediate visual feedback as to the effect of a joint: a hinge joint node with a certain origin or angle extent input will immediately transform the child geometry in the GUI according to those inputs, which prevents erroneous values and can be used to debug any arithmetic expressions responsible for calculating the joint parameters. Third, this standardized implementation of joints allows us to add logic to save all joint metadata in a standard format for later use.
@@
Our joint nodes support the creation of many articulated structure types such as jointing multiple meshes together in a chain, jointing multiple meshes to one parent, and connecting two geometries via multiple joints. We also create an additional custom node group that supports duplicating jointed bodies at defined points. This can be effective for procedurally duplicating articulations such as burners on a cooktop or shelves in a dishwasher without having to manually define each joint. Examples of all these cases can be found in Appendix Sec. A.2. Our joint nodes also optionally take as arguments semantic labels for the joint and the parent and child bodies. We develop an additional custom node that users can directly inject into their graph for more fine-grained labeling of individual parts. Finally, we provide a utility script that checks if any two rigid parts of the asset penetrate each other within a defined range of motion. This helps users to adapt joint configurations and parts’ geometries to avoid self-collision and penetration problems.
3.2 Exporting Assets to Simulation
Converting procedural articulated assets from Blender to a simulation-ready format involves three steps, all of which are automated. First, given the user’s articulated object in Blender, we use the Blender Python API to parse and update the asset’s node graph. During this step, we inject Blender-provided Store Named Attribute nodes into the original graph. These nodes assign values for attributes to different parts of the geometry that will be beneficial later during exporting. We also create a kinematic blueprint for the object class (inspired by [41, 42]) that abstracts away the full procedural graph while preserving the essential details required for articulation. We note that a kinematic blueprint is generated for an asset category, not an individual instance.
@@ 3.2 Exporting Assets to Simulation
Second, after generating a blueprint, we use Infinigen’s [38] transpiler to convert the updated node graph to Python code. Users can edit this code to customize asset parameter distributions or further tune node graph structures. Thus, our tool gives users fine-grained control over generated assets.
@@
The final step is to export the asset into a simulation-compatible format. Using the transpiled code, we sample random parameters and construct a new instance of the asset. We then pass this geometry along with its category’s corresponding kinematic blueprint to an exporter for either one of the URDF, USD, or MJCF file formats. The exporter recursively traverses the kinematic blueprint to build the final simulation-ready asset including all joints, rigid geometries, and semantic metadata. During this process, users can generate convex-decomposed collision meshes with third-party tools like CoACD [43].
Asset Category Discrete Variations Continous Variations
Door Door count: 1–2; Handle types (5): lever, knob, pullbar, crashbar, no handle; Door types (3): louver, glass, panel; Hinge sides (2): left, right. 39 dimensions
Fridge Door count: 1–2; External drawer count: 0–2+; Internal shelf count: 0–3+; Internal drawer count: 0–2+; Number of shelves per door (4): 0–4. 32 dimensions
Dishwasher Rack count: 0–3+; Button count: 0–6+; Button types (2): square, circle; Handle types (2): square, circle; Handle curvature (2): curved, square. 13 dimensions
Lamp Arm segments: 0–3; Type of joint for segment (2): rotational, translational; Switch types (4): rocker, twist, push, pull‐string; Switch location: base/head (n/a for pull‐string). 29 dimensions
Toaster Number of slots: 1–3+; Body protrusion axis (2): x‐axis, y‐axis; Levers per slot: 1–2; Types of levers (5): cylinder, half‐cylinder, ellipsoid, sphere, flat; Buttons per lever: 0–3+. 14 dimensions
Table 1: Procedural articulated object generators can densely cover a cartesian-product of possible subcomponents and continous parameters. For each object, we show the list of possible discrete variations (e.g. swappable sub-components) and calculate the total number of combinations of these components. We do not count material randomization. Then, we list the number of independent continous parameters (see Appendix Sec. A.3 for names), summed over subcomponents, which specifies the dimensionality of the continous parameter space. Combining discrete and continuous variations, and assuming 3 values per continous parameter, our system can generate between $10^{6}$ to $10^{20}$ unique assets per category.
4 Procedural Articulated Asset Generators
Using our custom joint and utility nodes, users can create a variety of physically accurate, kinematically diverse, photorealistic assets. We demonstrate the capabilities of Infinigen-Sim by developing procedural generators for five categories of articulated assets including doors, toasters, fridges, dishwashers, and lamps. These articulated procedural generators were created in 1-7 hours per object class by novice and intermediate Blender users, mostly dominated by time to create the static procedural asset.
Examples of assets sampled from these generators can be seen in Fig. 1. Table 1 showcases the variations we consider for each asset category. Similar assets in PartNet-Mobility [18] are quite limited in quantity.
@@ 4 Procedural Articulated Asset Generators
Figure 3: Comparison of asset parameter distributions. The top row shows the default distribution used for handle length and handle-to-hinge distance. The bottom row shows an expanded distribution where these parameters are varied over a wider range.
@@
All assets created in this way have many continously varying input parameters, which can either be manually set or randomized according to a distribution. We show a case study for just doors with lever handles in Fig. 3. To open a door with a handle, two parameters are important for manipulation: the length of the handle and the distance between the handle hinge and the door hinge. The former defines grasp poses and the motion of handle rotation. The latter determines the trajectory required to successfully rotate the door. Infinigen-Sim allows users to define widely distributed combinations of these two parameters with diverse handle geometries. Users can tailor distributions of parameters to their specific needs. Such properties cannot be satisfied with existing articulated assets [18, 19] or modeling tools [29, 33].
5 Experiments
We demonstrate that assets created with Infinigen-Sim help downstream vision and robot learning tasks, including movable part segmentation, RL generalization, and sim-to-real transfer.
5.1 Movable Part Segmentation
Movable part segmentation [44, 18] is an important vision task for embodied agents. Given an RGB image of an object, the model must segment each articulated part. We focus on five categories of assets including doors, toasters, refrigerators, dishwashers, and lamps. First, we generate 250 diverse assets per category using Infingen-Sim. Then, we follow a similar setup as Xiang et al. [18] to generate image datasets for both PartNet-Mobility and Infinigen-Sim assets. We split the PartNet dataset such that images rendered from 75% of the assets are used for training and the remaining 25% are used for evaluation. All models are evaluated on images of unseen PartNet assets only. We finetune a pretrained Mask R-CNN model with a ResNet-50-FPN backbone, as in [45], for three datasets: 1) P15k, approximately 15k images of PartNet assets only, 2) P30k, approximately 30k images of PartNet assets only, and 3) P15k+I, a combination of approximately 15k images of PartNet assets and 15k images of Infinigen-Sim assets. Please refer to Appendix A.4 for more details on data generation and model training.
@@ 5.1 Movable Part Segmentation
Table 2: Movable part segmentation performance per articulated part type. We train a model for 50 epochs and report the evaluation mAP scores averaged across 3 seeds. Using Infinigen-Sim assets results in stronger model generalization for relatively smaller articulated parts, e.g. door handles, toaster levers, dishwasher buttons, and lamps switches.
Through our experiments, we observe that doubling the size of the original dataset (P30k) yields only marginal improvements (see Appendix A.5 for overall results). This suggests the model may not be learning new features that improve generalization to unseen assets despite scaling the training images from PartNet assets alone. In contrast, adding the same number of images using Infinigen-Sim assets yields a larger performance gain. As shown in Table 2, this improvement is more significant across smaller articulated parts such as door handles, toaster levers, dishwasher buttons, and lamp switches. This suggests a core benefit of Infinigen-Sim. Namely, our tool allows users to procedurally scale articulated parts that require a greater level of detail. Achieving this level of granularity is labor-intensive for manually annotated articulation datasets.
5.2 Reinforcement Learning Generalization
Figure 4: Top row: RL environment setup in Maniskill3 [2]. From left to right: Push Down Toaster Lever, Push Door with Handle, Open Fridge Door. Bottom row: learning curve on the testing PartNet instances for all three tasks. Each curve shows the mean and standard deviation of 4 runs.
We demonstrate that Infinigen-Sim assets help reinforcement learning robot policies generalize to novel object instances. We evaluate on the following tasks.

### v3

# Procedural Generation of Articulated Simulation-Ready Assets

@@ Abstract
We introduce Infinigen-Articulated, a toolkit for generating realistic, procedurally generated articulated assets for robotics simulation. We include procedural generators for 18 common articulated object categories along with high-level utilities for use creating custom articulated assets in Blender. We also provide an export pipeline to integrate the resulting assets along with their physical properties into common robotics simulators. Experiments demonstrate that assets sampled from these generators are effective for movable object segmentation, training generalizable reinforcement learning policies, and sim-to-real transfer of imitation learning policies.
Figure 1: Procedurally generated assets for 18 common articulated object categories.
I INTRODUCTION
@@
Interacting with articulated objects is an essential step in many important applications for robotics. From using hinged fridges and dishwashers, pushing buttons, opening drawers, to traversing doors with various handles, robots must manipulate these articulated objects to perform useful real-world tasks.
@@
Learning in physical simulation [1, 2, 3, 4, 5, 6] has proven a useful strategy for many robotics tasks [7, 8, 9, 10, 11, 12, 13]. To best use this strategy for articulated object manipulation, we require large-scale datasets of articulated assets with as diverse geometry, joints, and visual appearance as possible. Articulated assets are challenging to acquire, as objects curated from the internet [14, 15] or 3D-scanning [16] are typically static and lack joint annotations. Widely used articulated object datasets instead use labor-intensive human annotation [17, 18, 19, 20, 21] for joint position and axes, which limits the quantity, quality, and diversity of unique geometries and articulation annotations.
To this end, we propose Infinigen-Articulated, a tool which enables the creation of high-quality, diverse articulated assets via procedural generation. Our tool has many advantages:
  1. 1.
@@
    Unlimited kinematic variation: Procedural generators sample assets with dense coverage of task-relevant parameters that influence a robot’s trajectory, e.g. the handle-to-hinge distance of a door. Continuous variation also applies to geometry detail (e.g. edge bevel radii) and tolerances (e.g. toaster slot width), which all serve to create dense coverage and diverse training data.
  2. 2.
    High-quality assets: Joint locations, axes, and annotations can be verified for correctness unlike previous efforts which rely on human annotation or generative methods. Our assets also utilize physics-based rendering materials present in Infinigen [22, 23] along with material physical properties such as friction and density.
  3. 3.
@@
    Combinatorial coverage: Procedural generators can randomly swap sub-parts to use different procedural generators (e.g. changes in handle style: knobs, levers, crash-bars). We can also easily express variable quantities of articulated parts (e.g. 1-5 fridge drawers), including very small components (e.g. tiny buttons on a toaster), which is challenging for annotators or image-based capture.
@@
Infinigen-Articulated builds on 3D-graphics and procedural generation tools from Blender [24] and Infinigen Indoors [23] by adding tools and generators designed for articulated object generation. Specifically, we augment Blender’s Geometry Nodes system with new modules (known as node-groups) which implement hinged and sliding joints for procedural assets. These node-groups ensure that joints are represented in a unified format for use by downstream tools. We provide export code which converts assets created using these tools into high quality assets for simulation in popular formats such as URDF, USD, and MJCF.
Experiments show that objects from procedural generators created with Infinigen-Articulated improve upon baselines in multiple settings for vision and robot learning. We use Infinigen-Articulated to create articulated procedural generators for 18 object categories. We hope Infinigen-Articulated will provide an ever-growing library of open-source articulated object generators as to enable more robust robot learning across many tasks using articulated objects.
II Related Work
@@
Asset Datasets for Robot Simulation. Although works have collected large-scale static 3D asset datasets [15, 14], interactive and articulated 3D datasets [17, 25, 26, 27, 21] remain limited in both quantity and quality. Sapien [17] collected 2k+ assets from PartNet [25] and manually annotated the articulated joints. Others [18, 19, 28] scanned real-world objects to curate separate articulated object datasets. However, issues like inaccurate joint annotations and a lack of kinematic diversity remain. Infinigen-Articulated instead adopts a new approach (procedural generation) which has precise ground truth and is highly scalable, only requiring a new procedural generator for each object class, with no human time cost for each unique object geometry or articulation configuration.
@@
Articulated Object Generation. Recent work has exploited AI techniques to augment and diversify static asset datasets [15, 29] to create articulated objects [30, 31]. However, these systems lack controllability of asset geometries and do not guarantee correct joint configurations [32]. Prior work [33, 30, 34, 35, 36, 37, 31, 38] has also generated articulated assets with real-to-sim pipelines. Paris [35] learns part-based neural radiance fields to reconstruct articulated assets. Real2Code [34] and URDFormer [36] constructs URDFs with RGB images as input, using a mixture of off-the-shelf vision models and LLMs. That being said, real-to-sim approaches often fail to construct articulated assets with high-quality. Additionally, these methods fall short on capturing occluded articulations and tend to produce inaccurate geometries and joint annotations [32]. Inherited from Infinigen-Indoors [23], our procedural generators model each individual part from scratch including their joint properties. This means users can also easily create high quality articulated bodies that may be occluded by outer surfaces (e.g. racks in a dishwasher or drawers in a refrigerator).
@@
Procedural Generation. Many prior works create procedural training data focused on indoor environments [4, 22], nature [22], and cities [39] but do not create procedural articulated objects. More similar to our work is [40], which introduces a scene generator including some procedural articulated assets made in Python. However, their assets are composed mainly of mesh primitives (cubes/cylinders), do not contain multi-level articulations (e.g. pivoting handle on a pivoting door), and is extensible only by programmers. Our system is constructed using Blender, has detailed geometry and materials, and is extensible by both programmers and anyone familiar with Blender.
III Infinigen-Articulated
Infinigen-Articulated consists of a variety of procedural generators for common articulated object categories along with utilities designed to aid in creating custom generators. We also develop export code which enables articulated objects to be used in robotics simulators. The overall workflow is visualized in Fig. 2.
III-A Articulated Procedural Generator Tools
@@
Figure 2: We first create procedural assets using Blender’s Geometry Nodes feature. These assets are then transpiled, and we have the ability to define custom distributions for the procedural parameters (including materials) along with its dynamics properties. We sample parameters to generate the asset and pass both the asset and its kinematic tree to the exporter, which recursively builds a simulation-ready articulated object.
@@
Infinigen-Articulated leverages Blender’s procedural modeling tools [24], namely Geometry Nodes and Shader Nodes, which are an artist-friendly, GUI-based, domain specific language for procedural generation. These tools allow users to create meshes and materials by composing nodes representing primitives, geometric transformations, and scalar vector arithmetic in a directed acyclic graph. These node-based tools are general purpose and can represent essentially all desirable object geometries.
@@
To design articulated assets within this system, we create two new nodes that implement revolute and prismatic joints. These nodes accept two incoming geometries (defined via prior nodes) and introduce a parent-child relationship annotated with the relevant joint type. Each joint node also allows users to set the pivot position, axis, and joint range, either as fixed values or through parameters passed from other nodes. For the provided generators, we leverage Blender’s built-in math nodes to compute these values directly from the asset’s geometry within the node graph.
@@
These joint nodes have four purposes. First, they speed up the procedural generator creation process since they implement a commonly used operation which no longer needs to be re-implemented for each object that uses a hinge or sliding joint. Second, they provide immediate visual feedback as to the effect of a joint: a hinge joint node with a certain origin or angle extent input will immediately transform the child geometry in the GUI according to those inputs, which prevents erroneous values and can be used to debug any arithmetic expressions responsible for calculating the joint parameters. Third, this standardized implementation of joints allows us to add logic to save all joint metadata in a standard format for later use. Fourth, these nodes can store semantic information, ensuring consistent, fine-grained joint and geometry naming across all procedurally generated assets within a class.
@@
Our joint nodes support the creation of many articulated structure types such as jointing rigid bodies in a sequence, jointing multiple bodies to a single parent, and adding multiple degrees of freedom between two bodies. We also create an additional custom node group that supports duplicating jointed bodies at defined points. This can be effective for procedurally duplicating articulations such as burners on a cooktop or shelves in a dishwasher without having to manually define each joint. Finally, we provide a utility script that checks if any two rigid parts of the asset penetrate each other within a defined range of motion. This helps users designing custom articulated assets adjust joint configurations and part geometries to prevent self-collisions and penetration issues.
III-B Exporting Assets to Simulation
@@
Converting instances of procedurally articulated assets from Blender to a simulation-ready format involves three steps, all of which are automated. First, we use Infinigen’s [22] transpiler to convert the procedural asset’s node graph to Python code. Users can edit this code to customize asset parameter distributions or further tune node graph structures. Thus, our tool gives users fine-grained control over generated assets. It also allows users to create dynamic graphs that may change based on the procedural parameters.
Second, we sample random values for each procedural parameter and spawn an instance of the asset, along with its node graph, in Blender. We then use the Blender Python API to parse and update the asset’s node graph. During this step, we inject into the graph Blender-provided nodes which assign values for attributes to different parts of the geometry. These attributes enable us to construct a kinematic tree for the asset instance (inspired by [41, 42]) that abstracts away the full procedural graph while preserving the essential details required for articulation.
@@
The final step is to export the asset into a simulation-compatible format. We pass the asset instance along with its corresponding kinematic tree to a native exporter for either one of the URDF, USD, and MJCF file formats. To construct the asset, the exporter recursively traverses the kinematic tree depth-first. To get individual rigid bodies, the exporter queries the asset instance with attributes and their values based on the current path in the tree. We then extract the part of the asset that matches this query and save it as an individual mesh. Users can also generate convex-decomposed collision meshes with third-party tools like CoACD [43]. During this process, we use the kinematic tree to build the final simulation-ready asset including all joints, rigid geometries, and semantic metadata.
IV Articulated Asset Procedural Generators
Using Infinigen-Articulated, we create procedural generators for 18 categories of common articulated assets as shown in Fig. 1. Each procedural generator includes distributions for joint dynamics parameters (i.e. stiffness and damping) and produces kinematically diverse, photorealistic articulated assets. A list of articulated parts is provided in VII-A.
@@ IV Articulated Asset Procedural Generators
Figure 3: Comparison of asset parameter distributions. The left column shows the default distribution used for handle length and handle-to-hinge distance. The right column shows an expanded distribution where these parameters are varied over a wider range.
@@
All assets created in this way have many continuously varying input parameters, which can either be manually set or randomized according to a distribution. We show a case study for doors with lever handles in Fig. 3. To open a door with a handle, two parameters are critical for manipulation: the length of the handle and the distance between the handle hinge and the door hinge. The former defines grasp poses and the motion of handle rotation. The latter determines the trajectory required to successfully rotate the door. Infinigen-Articulated allows users to define widely distributed combinations of these two parameters with diverse handle geometries. Users can tailor distributions of parameters to their specific needs. Such properties cannot be satisfied with existing articulated assets [17, 18] or modeling tools [30, 34].
Another key advantage of procedurally generating articulated assets is the ability to scale small, detailed components. These parts are particularly important to produce robust policies for tasks such as opening a door with a lock or pulling a thin, potentially occluded, chain to turn on a lamp. Our provided procedural generators include several detailed parts, each with their own procedural parameters. This provides users the ability to scale assets with high levels of detail and accuracy (i.e. correct joint position, axis, and range) while incurring little overhead. Examples of such articulated parts can be seen in Fig. 4. In contrast, manually annotating and articulating detailed parts of an asset is time-consuming and is much harder to scale accurately.
Figure 4: Examples of articulations for small, but essential parts (rows 1 and 2), occluded objects (rows 3 and 4), and double-jointed objects (row 5).
@@
TABLE I: Movable part segmentation performance per articulated part type. We train a model for 50 epochs and report the evaluation mAP scores averaged across 3 seeds. Using Infinigen-Articulated assets results in stronger model generalization for relatively smaller articulated parts, e.g. door handles, toaster levers, dishwasher buttons, and lamp switches.
V Experiments
We demonstrate Infinigen-Articulated’s potential on downstream vision and robot learning tasks, including movable part segmentation, RL generalization, and sim-to-real transfer.
V-A Movable Part Segmentation
Movable part segmentation [44, 17] is an important vision task for embodied agents. Given an RGB image of an object, the model must segment each articulated part. We focus on five categories of assets including doors, toasters, refrigerators [^1] The refrigerators used in the movable part segmentation experiments have since been updated (Fig. 1) with more diverse handle types, more realistic materials, and fewer vertices., dishwashers, and lamps. First, we generate 250 diverse assets per category using Infingen-Sim. Then, we follow a similar setup as [17] to generate image datasets for both PartNet-Mobility and Infinigen-Articulated assets. We split the PartNet dataset such that images rendered from 75% of the assets are used for training and the remaining 25% are used for evaluation. All models are evaluated on images of unseen PartNet assets only. We finetune a pretrained Mask R-CNN model with a ResNet-50-FPN backbone, as in [45], for three datasets: 1) P15k, approximately 15k images of PartNet assets only, 2) P30k, approximately 30k images of PartNet assets only, and 3) P15k+I, a combination of approximately 15k images of PartNet assets and 15k images of Infinigen-Articulated assets.
As shown in Table I, we observe that doubling the size of the original dataset (P30k) yields only marginal improvements, and in some cases decreases performance. This suggests a saturation in the model’s ability to learn new features that improve generalization to unseen assets despite scaling the training images from the original PartNet dataset alone. In contrast, we see a significant improvement when co-training with images of Infinigen-Articulated assets across smaller articulated parts such as door handles, toaster levers, dishwasher buttons, and lamp switches. This highlights a core benefit of Infinigen-Articulated. Namely, our tool allows users to procedurally scale articulated parts that require a greater level of detail. Achieving this level of granularity is labor-intensive for manually annotated articulation datasets. Despite our assets being out of distribution compared to the evaluation set, we still see improvements for detailed parts.
V-B RL Generalization
We demonstrate that Infinigen-Articulated assets help reinforcement learning policies generalize to novel object instances. For each task, we train three policies, namely one trained with only Infinigen-Articulated assets, one trained with only Partnet-Mobility assets, and one trained on a combination of both. We evaluate on the following tasks: